# Where Do We Stand with Implicit Neural Representations? A Technical and Performance Survey

### v1

@@ 1 Introduction
Knowledge representation (Brachman, 2004) is a foundation in computational fields, playing a critical role in enabling systems to efficiently model, interpret, and manipulate information across various domains. Deep neural networks have demonstrated a powerful capacity for learning robust knowledge representation from data, and have become the predominant tools for addressing complex tasks in areas like computer vision (LeCun et al., 2015). The significance of effective knowledge representation extends beyond traditional methods, as it directly influences the performance and scalability of systems when handling diverse types of information such as images, video, and audio, whether in 1D, 2D, or 3D formats. Conventional approaches to encoding input signals typically rely on explicit discretisation, where the input space is segmented into distinct elements, such as point clouds (Achlioptas et al., 2018; Fan et al., 2017), voxel grids (Gadelha et al., 2017; Liao et al., 2018; Stutz & Geiger, 2018; Jimenez Rezende et al., 2016), and meshes (Kanazawa et al., 2018; Ranjan et al., 2018; Wang et al., 2018). While these methods can achieve adequate results, they present significant challenges when dealing with high-dimensional data (Mescheder et al., 2019). The computational cost of discretisation rises sharply with increasing dimensionality, making it inefficient, particularly for complex or irregular spaces. Moreover, traditional discretisation methods tend to require substantial memory, posing limitations for large-scale applications. Implicit Neural Representations (INRs) offer a promising alternative by using continuous functions to represent data, addressing many of the limitations of explicit methods, such as memory inefficiency and the high computational cost associated with discretisation.
@@
The rise and continued development of INR have recently emerged as a new way to learn representation efficiently. Implicit representations differ from explicit (or discrete) representations by encoding information as a continuous generator function, which maps input coordinates to corresponding values within the defined input space, rather than directly storing feature or signal values. Consequently, there has been significant interest in utilising these networks as implicit functions, with notable success (Mildenhall et al., 2021; Mescheder et al., 2019; Xie et al., 2022). Specifically, Multi-Layer Perceptrons (MLPs) are trained to parameterise signals by taking input coordinates by a mapping technique that projects the network input into a higher-dimensional space. It predicts the associated data values. In this framework, the MLP functions as an Implicit Neural Representation, encoding the signal’s information within its weights. For example, when applied to image data, pixel coordinates are fed into the MLP, which generates the corresponding RGB values, effectively learning a continuous, high-resolution representation of the image. However, the classic use of the ReLU activation function often resulted in suboptimal performance across many applications. To address this issue, reparametrised learning techniques (Rahaman et al., 2019) have been first used to adjust the weights and mitigate bias, further enhancing the network’s performance. (Fathony et al., 2020) presented a new architecture where the output of each layer is multiplied by a Gabor wavelet. Further researchers have introduced various activation functions. These include periodic sinusoidal functions (Sitzmann et al., 2020), time-frequency localised Gabor wavelets (Saragadam et al., 2023), Gaussian functions (Ramasinghe & Lucey, 2022), and the FINER network (Liu et al., 2024). Additionally, Trident (Shen et al., 2023) is a network that integrates both positional encoding and a carefully chosen activation function.
@@
Implicit neural functions have been further adjusted in various tasks, including image generation (Reddy et al., 2021), super-resolution (Wu et al., 2021; Chen et al., 2021), 3D object reconstruction (Chabra et al., 2020; Mescheder et al., 2019; Mildenhall et al., 2021), and modelling of complex signals (Xu et al., 2022). The use of multi-layer perceptrons (MLPs) for image and shape parameterisation provides distinct advantages. First, MLPs are resolution-independent as they operate within a continuous domain, enabling them to generate values for coordinates beyond pixel- or voxel-based grids. Thus, it improves performance in vision tasks. Second, their memory requirements are not constrained by signal resolution, allowing more memory-efficient representations compared to traditional grid or voxel methods (Huang et al., 2021; Park et al., 2019). The memory demand scales according to the complexity of the signal rather than the resolution. Additionally, MLPs address the limitations of locality biases often found in convolutional neural networks (CNNs), which can hinder generalisation (Chen & Zhang, 2019). Finally, MLP-based models are fully differentiable, offering adaptability across various applications (Zhang et al., 2024; Liu et al., 2020). Their weights can be optimised using gradient-based techniques, providing the flexibility needed for diverse tasks (Zhang et al., 2023; Xie et al., 2022; Tancik et al., 2020; Cheng et al., 2023).
@@ 2 Where Do We Stand with Implicit Neural Representations?
$F(x,a(x),\nabla_{x}a(x),\nabla^{2}_{x}a(x),...)=0$ (1)
Our function of interest could be simply be the image, mapping the coordinates to their pixel values. To do that, we parametrise $a(x)$ as an MLP network with N layer, where the output of each layer is given by:
@@
where $\sigma$ is the activation function, $W_{i}\in\mathbb{R}^{n_{i}\dot{n}_{i-1}}$ are the weights, and $b_{i}\in\mathbb{R}^{n_{i}}$ the biases of the i-th layer. $x_{0}$ is the input coordinate (i.e., pixel coordinate for an image), and the final output reads: $y_{M}=W_{N}y_{N-1}+b_{N}$
@@ 2.1 Taxonomy INR: (a) Activation Function
To begin, we present our first category in our taxonomy and review the existing literature. One of the critical challenges in Implicit Neural Representations (INRs) is overcoming the bias towards learning low-frequency components, commonly observed in neural networks that use traditional activation functions such as ReLU (Hanin & Rolnick, 2019). The ReLU function, while effective in many settings, has a discontinuous nature that limits its ability to capture higher-frequency signals, making it suboptimal for tasks that require detailed or high-resolution data representation. Furthermore, alternative activation functions such as sigmoid or tanh, while continuous, are not expressive enough to adequately capture the full range of frequency components necessary for complex tasks.
@@
$\sigma(x)=sin(\omega x)$ (3)
where $\omega$ is a scaling hyperparameter chosen tailored for each task to modulate the low-frequency bias. This activation function stands out due to its periodic nature, enabling it to model a broad spectrum of frequencies effectively. The periodic properties of Siren allow the network to represent high-frequency variations with ease, making it particularly suitable for tasks requiring intricate details, such as image and audio reconstruction, and solving complex differential equations. One key benefit of the sine-based activation function is its smoothness, which not only facilitates stable gradient flow during training but also allows for explicit calculation of derivatives. This capability makes Siren highly effective for applications involving inverse problems, where precise derivative computation is critical. Additionally, the smooth, continuous representation that Siren offers enables it to generalise well across varying resolutions and achieve high-quality, resolution-independent outputs. The authors propose a specialised weight initialisation scheme to maintain the distribution of activations throughout the network layers, which ensures efficient convergence and avoids issues like vanishing or exploding gradients. This initialisation technique further enhances Siren’s capacity to capture fine-grained details while remaining memory efficient.
@@
Wire (Saragadam et al., 2023) The authors propose using the Gabor wavelet as an activation function, defined as:
$\sigma(x)=\exp(-(sx)^{2}+i\omega x)$ (5)
The Gabor wavelet is effective in minimising the product of its standard deviations in both the time and frequency domains, allowing it to represent features across both space and frequency. This dual localisation enables the function to capture fine-grained details and frequency components more effectively. As a result, the Gabor wavelet combines the advantages of both the Siren network, which excels in high-frequency representations, and the Gaussian network, known for its smooth, spatial localisation.
While the Gabor wavelet offers strong representational capabilities, a potential drawback is its relatively higher computational cost compared to simpler activation functions. The need to process both spatial and frequency components simultaneously can lead to increased complexity in network training and inference, making it less efficient in scenarios where computational resources or speed are critical. In addition, the Gabor wavelet’s ability to capture fine-grained details across both space and frequency, while advantageous, may also increase the risk of overfitting, particularly when dealing with noisy or sparse data. The model may become overly sensitive to minor variations in the input, making it less generalisable to unseen data.
@@
$\sigma(x)=\tanh(\beta\sin(x))$ (6)
@@
Sinc Function (Saratchandran et al., 2024) Motivated by the classic Shannon sampling theorem and drawing similarities between INR and sampling, the authors in (Saratchandran et al., 2024) proposed the sinus cardinal as an activation function, which form is give by:
$\sigma(x)=\text{sinc}(\omega x)$ (7)
The sinus cardinal is localised in the space domain, but not in the frequency domain. It also forms a generating system, meaning the set of its translates $\{x\to\text{sinc}(x-k)|k\in\mathbb{Z}\}$ can be used for sampling, enabling the network to approximate signals and represent high-order features effectively. The Sinc activation function, motivated by sampling theory, offers several distinct advantages. One of its primary strengths is its localisation in the space domain, which makes it particularly effective at representing spatial information. Additionally, the sinus cardinal function can generate a set of translates that approximate signals, making it well-suited for tasks that involve high-order feature representation. Its connection to sampling theory provides a theoretical foundation for reconstructing signals, which enables accurate approximation of continuous data.
However, a key limitation of the Sinc function is its lack of localisation in the frequency domain. This means it might struggle to represent or capture high-frequency variations as effectively as other functions designed with periodicity or high-frequency adaptivity. As a result, while the Sinc function excels in certain applications where spatial representation is paramount, it may not be the best choice for tasks that require precise handling of complex, high-frequency features.
@@
Finer (Liu et al., 2024) The Siren model struggles to represent a broad spectrum of frequencies due to its reliance on a fixed scaling value. This limitation restricts the range of reconstructed frequencies, which may be insufficient, particularly when the frequency distribution of the signal is unknown or varies significantly. To overcome this issue, the authors of that (Liu et al., 2024) proposed an activation function given by:
$\sigma(x)=\sin(\omega(|x|+1)x)$ (8)
This activation function is both smooth and periodic, inheriting all the properties of the sine function. In addition, the scaling parameter $\omega$ varies dynamically across the network based on the bias of different nodes. Higher bias values correspond to rapidly varying functions that can capture high frequencies, while lower bias values correspond to slower variations, effectively representing lower frequencies. This dynamic scaling provides the Finer approach with greater flexibility to represent a broader range of frequencies. Specifically, nodes with higher bias can capture high-frequency details, while nodes with lower bias are better suited for representing lower-frequency information. This adaptability offers Finer a major advantage over fixed-scaling models like Siren, as it is able to represent a much broader range of frequencies, particularly when the frequency distribution is variable or unknown. Accordingly, the authors to initialise the bias coefficients as:
$b\sim U(-k,k)$ (9)
@@ 2.3 Taxonomy INR: (c) Activation Function + Position Encoding
Trident (Shen et al., 2023) Trident is a mix between the two approaches of using an activation function, and applying a Fourier mapping. The first layer of the network is expressed as:
@@
$cos(x)^{n}=\frac{1}{2^{n}}\binom{n}{\frac{n}{2}}+\frac{2}{2^{n}}\sum_{k=0}^{n}{\frac{n}{2}}\binom{n}{k}cos((n-2k)x)$ (13)
@@ 2.4 Taxonomy INR: (d) Network
$\sigma(x;\theta^{i})=\sin(\omega^{i}x+\phi^{i})$ (15)
where the frequency $\omega^{i}$ and phase shift $\phi^{i}$ allow the model to capture oscillatory patterns effectively. The second is the Gabor wavelet, given by:
$\sigma((x;\theta^{i})=\exp(-\gamma^{i}\|x-\mu^{i}\|_{2}^{2})\sin(\omega^{i}x+\phi^{i}),$ (16)
which introduces both spatial and frequency localisation. The combination of these functions enables Mfn to model a wide range of data patterns by approximating signals as a mixture of sinusoidal and wavelet components. One of the key advantages of Mfn is their ability to represent complex, high-order details that are often challenging for standard architectures like Siren or Fourier Feature Networks (Fourier Features). By using multiplicative interactions and flexible nonlinearities, Mfn can capture intricate dependencies within the input data. This makes them particularly effective in applications requiring fine-grained representation and high-frequency details, such as audio synthesis, image reconstruction, and other tasks involving implicit neural representations.
@@
Fr–Fourier Reparameterised Training (Shi et al., 2024) Another approach to bypass the low frequency bias is to use an appropriate weight reprametisation of weights. The idea is to reparametrise each row of the weight matrix of the $i$-th layer $W_{i}$ as a sum of Fourier bases. The output of each layer is given by:
@@
$B_{i}^{k,l}=cos(w_{k}z_{l}+\phi_{k})$ (18)
for $i=1,...,M$ and $j=1,...,n_{out}$. Where $B_{i}^{k,l}$ are the entries of $B_{i}$, $(z_{l})$ is the sampling position sequence and $M$ is the number of Fourier basis considered.
The authors proposed to choose $P$ phases as an array of phase shifts, which are evenly distributed over the interval from $[0,2\pi]$, and to take $2F$ frequencies, consisting of a low frequency basis $\{\frac{1}{F},\frac{2}{F},...,1\}$ and a high frequency basis $\{1,2,...,F\}$, which gives a total of $M=2FP$ bases. $F$ and $P$ are hyper parameters and should be chosen for each task independently.

### v2

@@ 1 Introduction
Knowledge representation (Brachman, 2004) is a foundation in computational fields, playing a critical role in enabling systems to efficiently model, interpret, and manipulate information across various domains. Deep neural networks have demonstrated a powerful capacity for learning robust knowledge representation from data, and have become the predominant tools for addressing complex tasks in areas like computer vision (LeCun et al., 2015). The significance of effective knowledge representation extends beyond traditional methods, as it directly influences the performance and scalability of systems when handling diverse types of information such as images, video, and audio, whether in 1D, 2D, or 3D formats. Conventional approaches to encoding input signals typically rely on explicit discretisation, where the input space is segmented into distinct elements, such as point clouds (Achlioptas et al., 2018; Fan et al., 2017), voxel grids (Gadelha et al., 2017; Liao et al., 2018; Stutz & Geiger, 2018; Jimenez Rezende et al., 2016), and meshes (Kanazawa et al., 2018; Ranjan et al., 2018; Wang et al., 2018). While these methods can achieve adequate results, they present significant challenges when dealing with high-dimensional data (Mescheder et al., 2019). The computational cost of discretisation rises sharply with increasing dimensionality, making it inefficient, particularly for complex or irregular spaces. Moreover, traditional discretisation methods tend to require substantial memory, posing limitations for large-scale applications. Recent works have shown efficient representation of complex data, for instance latent representation, as seen in autoencoders and variational autoencoders (VAEs) (Tschannen et al., 2018), which encode signals into low-dimensional latent spaces through the use of neural networks, and structured dictionary learning (Chen et al., 2018; Tasissa et al., 2023), which seeks to approximate signals as sparse linear combinations of learned basis elements. Implicit Neural Representations (INRs) offer a promising alternative by using continuous functions to represent data, addressing many of the limitations of explicit methods, such as memory inefficiency and the high computational cost associated with discretisation.
@@
The rise and continued development of INR have recently emerged as a new way to learn representation efficiently. Implicit representations differ from explicit (or discrete) representations by encoding information as a continuous generator function, which maps input coordinates to corresponding values within the defined input space, rather than directly storing feature or signal values. Consequently, there has been significant interest in utilising these networks as implicit functions, with notable success (Mildenhall et al., 2021; Mescheder et al., 2019; Xie et al., 2022). Specifically, Multi-Layer Perceptrons (MLPs) are trained to parameterise signals by taking coordinates as inputs, applying a mapping technique that projects this input into a higher-dimensional space, and predicting the associated data values at each coordinate. In this framework, the MLP functions as an Implicit Neural Representation, encoding the signal’s information within its weights. For example, when applied to image data, pixel coordinates are fed into the MLP, which generates the corresponding RGB values, effectively learning a continuous, high-resolution representation of the image. However, the classic use of the ReLU activation function often resulted in suboptimal performance across many applications. To address this issue, reparametrised learning techniques (Rahaman et al., 2019) have been first used to adjust the weights and mitigate bias, further enhancing the network’s performance. (Fathony et al., 2020) presented a new architecture where the output of each layer is multiplied by a Gabor wavelet. Further researchers have introduced various activation functions. These include periodic sinusoidal functions (Sitzmann et al., 2020), time-frequency localised Gabor wavelets (Saragadam et al., 2023), Gaussian functions (Ramasinghe & Lucey, 2022), and the Finer network (Liu et al., 2024). Additionally, Trident (Shen et al., 2023) is a network that integrates both positional encoding and a carefully chosen activation function.
@@
Implicit neural functions have been further adjusted in various tasks, including image generation (Reddy et al., 2021), super-resolution (Wu et al., 2021; Chen et al., 2021), 3D object reconstruction (Chabra et al., 2020; Mescheder et al., 2019; Mildenhall et al., 2021), and modelling of complex signals (Xu et al., 2022). Its application on modeling 3D scenes has been developed as Neural Radiance Field scene representations (NeRFs) (Mildenhall et al., 2021) allowing for photorealistic image generation from multiple viewpoints. The use of multi-layer perceptrons (MLPs) for image and shape parameterisation provides distinct advantages. First, MLPs are resolution-independent as they operate within a continuous domain, enabling them to generate values for coordinates beyond pixel- or voxel-based grids. Thus, it improves performance in vision tasks. Second, their memory requirements are not constrained by signal resolution, allowing more memory-efficient representations compared to traditional grid or voxel methods (Huang et al., 2021; Park et al., 2019). The memory demand scales according to the complexity of the signal rather than the resolution. Additionally, MLPs address the limitations of locality biases often found in convolutional neural networks (CNNs), which can hinder generalisation (Chen & Zhang, 2019). Finally, MLP-based models are fully differentiable, offering adaptability across various applications (Zhang et al., 2024; Liu et al., 2020). Their weights can be optimised using gradient-based techniques, providing the flexibility needed for diverse tasks (Zhang et al., 2023; Xie et al., 2022; Tancik et al., 2020; Cheng et al., 2023). Another body of work is that of (Shenouda et al., 2024). The authors presented insights that may complement the broader discussions in this survey, particularly in exploring the potential of ReLU-based architectures for INR tasks.
@@
Implicit Neural Representations (INRs) are increasingly intersecting with broader trends in Artificial Intelligence (AI). For instance, the adoption of INR in transformer architectures enables more expressive and scalable mechanisms for capturing complex data relationships (Chen & Wang, 2022). Another promising direction involves diffusion models that leveraging INR as guide (Hui et al., 2024) expanding their potential applications. Within this expanding landscape, INRs hold a unique position as compact, continuous representations of data, complementing other AI paradigms and offering a flexible tool for solving diverse problems in machine learning and beyond.
@@
Figure 3: Visualisation of the activation functions used in the method, categorised as (a) in Figure 1.
@@ 2 Where Do We Stand with Implicit Neural Representations?
$F(x,a(x),\nabla_{x}a(x),\nabla^{2}_{x}a(x),...)=0.$ (1)
Our function of interest could be simply be the image, mapping the coordinates to their pixel values. To do that, we parametrise $a(x)$ as an MLP network with N layers, where the output of each layer is given by:
@@
where $\sigma$ is the activation function, $W_{i}\in\mathbb{R}^{n_{i}\dot{n}_{i-1}}$ are the weights, and $b_{i}\in\mathbb{R}^{n_{i}}$ the biases of the i-th layer. $x_{0}$ is the input coordinate (i.e., pixel coordinate for an image), and the final output reads: $y_{M}=W_{N}y_{N-1}+b_{N}$.
@@ 2.1 Taxonomy INR: (a) Activation Function
To begin, we present our first category in our taxonomy and review the existing literature. One of the critical challenges in Implicit Neural Representations (INRs) is overcoming the bias towards learning low-frequency components, commonly observed in neural networks that use traditional activation functions such as ReLU (Hanin & Rolnick, 2019). The ReLU function, while effective in many settings, has a discontinuous nature that limits its ability to capture higher-frequency signals, making it suboptimal for tasks that require detailed or high-resolution data representation. However, recent work Shenouda et al. (2024) has demonstrated that by imposing specific constraints inspired by B-spline wavelets, ReLU-based networks can overcome this limitation, enabling them to approximate high-frequency components effectively while maintaining their computational efficiency. Furthermore, alternative activation functions such as sigmoid or tanh, while continuous, are not expressive enough to adequately capture the full range of frequency components necessary for complex tasks.
@@
$\sigma(x)=\sin(\omega x),$ (3)
where $\omega$ is a scaling hyperparameter chosen tailored for each task to modulate the low-frequency bias. This activation function stands out due to its periodic nature, enabling it to model a broad spectrum of frequencies effectively. The periodic properties of Siren allow the network to represent high-frequency variations with ease, making it particularly suitable for tasks requiring intricate details, such as image and audio reconstruction, and solving complex differential equations. One key benefit of the sine-based activation function is its smoothness, which not only facilitates stable gradient flow during training but also allows for explicit calculation of derivatives. This capability makes Siren highly effective for applications involving inverse problems, where precise derivative computation is critical. Additionally, the smooth, continuous representation that Siren offers enables it to generalise well across varying resolutions and achieve high-quality, resolution-independent outputs. The authors proposed a specialised weight initialisation scheme to maintain the distribution of activations throughout the network layers, which ensures efficient convergence and avoids issues like vanishing or exploding gradients. This initialisation technique further enhances Siren’s capacity to capture fine-grained details while remaining memory efficient.
@@
Wire (Saragadam et al., 2023) The authors proposed using the Gabor wavelet as an activation function, defined as:
$\sigma(x)=\exp(-(sx)^{2}+i\omega x),$ (5)
where the imaginary part is represented by the $i$-term. The Gabor wavelet is effective in minimising the product of its standard deviations in both the time and frequency domains, allowing it to represent features across both space and frequency. This dual localisation enables the function to capture fine-grained details and frequency components more effectively. As a result, the Gabor wavelet combines the advantages of both the Siren network, which excels in high-frequency representations, and the Gaussian network, known for its smooth, spatial localisation.
While the Gabor wavelet offers strong representational capabilities, a potential drawback is its relatively higher computational cost compared to simpler activation functions. The need to process both spatial and frequency components simultaneously can lead to increased complexity in network training and inference, making it less efficient in scenarios where computational resources or speed are critical. In addition, the Gabor wavelet’s ability to capture fine-grained details across both space and frequency, while advantageous, may also increase the risk of overfitting, particularly when dealing with noisy or sparse data. The model may become overly sensitive to minor variations in the input, making it less generalisable to unseen data. While most of the INR activation functions only require one hyperparameter, Wire requires two hyperparameters ($s$ and $\omega$), which in turn results in more computational effort to choose the appropriate hyperparameters.
@@
$\sigma(x)=\tanh(\beta\sin(x)),$ (6)
@@
Sinc Function (Saratchandran et al., 2024) Motivated by the classic Shannon sampling theorem and drawing similarities between INR and sampling, the authors in (Saratchandran et al., 2024) proposed the sinus cardinal as an activation function, which form is given by:
$\sigma(x)=\text{sinc}(\omega x).$ (7)
The Sinc is localised in the space domain, but not in the frequency domain. It also forms a generating system, meaning the set of its translates $\{x\to\text{sinc}(x-k)|k\in\mathbb{Z}\}$ can be used for sampling, enabling the network to approximate signals and represent high-order features effectively. The Sinc activation function, motivated by sampling theory, offers several distinct advantages. One of its primary strengths is its localisation in the space domain, which makes it particularly effective at representing spatial information. Additionally, the Sinc function can generate a set of translates that approximate signals, making it well-suited for tasks that involve high-order feature representation. Its connection to sampling theory provides a theoretical foundation for reconstructing signals, which enables accurate approximation of continuous data.
However, a key limitation of the Sinc is its lack of localisation in the frequency domain. This means it might struggle to represent or capture high-frequency variations as effectively as other functions designed with periodicity or high-frequency adaptivity. As a result, while the Sinc excels in certain applications where spatial representation is paramount, it may not be the best choice for tasks that require precise handling of complex, high-frequency features.
@@
Finer (Liu et al., 2024) The Siren model struggles to represent a broad spectrum of frequencies due to its reliance on a fixed scaling value. This limitation restricts the range of reconstructed frequencies, which may be insufficient, particularly when the frequency distribution of the signal is unknown or varies significantly. To overcome this issue, the authors of the work (Liu et al., 2024) proposed an activation function given by:
$\sigma(x)=\sin(\omega(|x|+1)x).$ (8)
This activation function is both smooth and periodic, inheriting all the properties of the sine function. In addition, the scaling parameter $\omega$ varies dynamically across the network based on the bias of different nodes. Higher bias values correspond to rapidly varying functions that can capture high frequencies, while lower bias values correspond to slower variations, effectively representing lower frequencies. This dynamic scaling provides the Finer approach with greater flexibility to represent a broader range of frequencies. Specifically, nodes with higher bias can capture high-frequency details, while nodes with lower bias are better suited for representing lower-frequency information. This adaptability offers Finer a major advantage over fixed-scaling models like Siren, as it is able to represent a much broader range of frequencies, particularly when the frequency distribution is variable or unknown. Accordingly, the authors proposed to initialise the bias coefficients as an uniform distribution:
$b\sim U(-k,k).$ (9)
@@ 2.3 Taxonomy INR: (c) Activation Function + Position Encoding
Trident (Shen et al., 2023) Trident is a mix between the two approaches of
using an activation function, and applying a Fourier mapping. The first layer of the network is expressed as:
@@
$\cos(x)^{n}=\frac{1}{2^{n}}\binom{n}{\frac{n}{2}}+\frac{2}{2^{n}}\sum_{k=0}^{n}{\frac{n}{2}}\binom{n}{k}\cos((n-2k)x).$ (13)
@@ 2.4 Taxonomy INR: (d) Network
$\sigma(x;\theta^{i})=\sin(\omega^{i}x+\phi^{i}),$ (15)
where in the $i$th layer, the parameter $\theta^{i}$ includes the frequency $\omega^{i}$ and phase shift $\phi^{i}$ allow the model to capture oscillatory patterns effectively. The second is the Gabor wavelet, given by:
$\sigma(x;\theta^{i})=\exp(-\gamma^{i}\|x-\mu^{i}\|_{2}^{2})\sin(\omega^{i}x+\phi^{i}),$ (16)
which introduces both spatial and frequency localisation with an additional parameter, the scale term $\gamma^{i}$. The combination of these functions enables Mfn to model a wide range of data patterns by approximating signals as a mixture of sinusoidal and wavelet components. One of the key advantages of Mfn is their ability to represent complex, high-order details that are often challenging for standard architectures like Siren or Fourier Feature Networks (Fourier Features). By using multiplicative interactions and flexible nonlinearities, Mfn can capture intricate dependencies within the input data. This makes them particularly effective in applications requiring fine-grained representation and high-frequency details, such as audio synthesis, image reconstruction, and other tasks involving implicit neural representations.
@@
INR Baseline Complexity Extra Complexity
Siren The sinusoidal operation is computed in constant time, with no additional complexity beyond $O(1)$.
Gauss The Gaussian function involves squaring the input, multiplying, and applying the exponential function, all of which are $O(1)$.
Wire Combines Gaussian $O(1)$ operations with cosine $O(1)$, resulting in no additional complexity beyond $O(1)$.
Hosc Combines sine $O(1)$ with hyperbolic tangent $O(1)$. The sharpness factor $\beta$ introduces no additional complexity.
Sinc Involves sine and division operations. The division introduces negligible overhead, remaining $O(1)$.
Finer $O(1)$ The sinusoidal operation with dynamic scaling based on $|x|$ introduces no additional asymptotic complexity, as absolute value and addition are $O(1)$.
Table 1: Computational Complexity of Activation Functions in INR Methods.
Fr–Fourier Reparameterised Training (Shi et al., 2024) Another approach to bypass the low-frequency bias is to use an appropriate weight reparametrisation. The idea is to reparametrise each row of the weight matrix of the $i$-th layer $W_{i}$ as a sum of Fourier bases. The output of each layer is given by:
@@
$B_{i}^{k,l}=\cos(w_{k}z_{l}+\phi_{k})$ (18)
for $i=1,...,M$ and $j=1,...,n_{out}$. We denote $B_{i}^{k,l}$ as the entries of $B_{i}$, $(z_{l})$ which is the sampling position sequence, and $M$ as the number of Fourier basis considered.
The authors proposed to choose $P$ phases as an array of phase shifts, which are evenly distributed over the interval from $[0,2\pi]$, and to take $2F$ frequencies, consisting of a low-frequency basis $\{\frac{1}{F},\frac{2}{F},...,1\}$ and a high frequency basis $\{1,2,...,F\}$, which gives a total of $M=2FP$ bases. $F$ and $P$ are hyper parameters and should be chosen for each task independently.
@@
2.5 On the Complexity of INRs
The Table 1 provides an overview of the computational complexity associated with various activation functions used in implicit neural representations (INRs). All the functions listed share a baseline computational complexity of $O(1)$, which means that evaluating these functions for a single input requires constant time. This characteristic is crucial for INR methods, as they often deal with high-dimensional data where computational efficiency is paramount. Each activation function in the table demonstrates $O(1)$ complexity, but they vary in the operations they involve and their ability to handle specific tasks. For instance, the Siren function relies solely on a sine operation, making it straightforward and efficient with no additional computational overhead. This simplicity makes it well-suited for representing periodic and oscillatory patterns in data. It can be proven from the results that it is the second fastest, 482 seconds, among all tested methods for the dosing task reported in Table 3. The Gaussian (Gauss) activation function, while also $O(1)$, requires squaring the input, multiplying it by a scalar, and applying an exponential operation. These operations remain computationally lightweight, with empirically reported 709 seconds for denoising task in Table 3, yet they provide the function with the ability to represent smooth and spatially localised features, which is especially useful for tasks involving smooth transitions. Wire takes this further by combining the Gaussian operation with a cosine function. Although this combination adds slight computational overhead compared to standalone functions like Siren or Gauss (as verified in Table 3 for the denoising task, 1485 seconds, which is double the computational time of Gauss and triple that of Siren), it still retains the $O(1)$ complexity. The dual spatial and frequency localisation capabilities provided by Wire justify this marginal increase in computational effort, making it versatile for representing more complex patterns. Hosc builds on a sine function by applying a hyperbolic tangent transformation. While this adds an extra step, both the sine and hyperbolic tangent functions are computed in constant time, maintaining $O(1)$ complexity. The flexibility introduced by the sharpness factor $\beta$ allows HOSC to adapt between high- and low-frequency signal representations without compromising computational efficiency. The Sinc, inspired by the sinus cardinal, involves both a sine computation and a division operation. Although the division introduces a minor computational overhead, it is negligible and does not affect the overall $O(1)$ complexity. Sinc’s strong spatial localisation capabilities make it particularly valuable for signal reconstruction tasks, though its frequency adaptability may be more limited than other functions.